### v2

# Quality Tagging, and Management of Synthetic Data

Abstract
The advancement of large language models (LLMs) is critically dependent on the availability of high-quality datasets for Supervised Fine-Tuning (SFT), alignment tasks like Direct Preference Optimization (DPO), etc. In this work, we present a comprehensive synthetic data generation framework, SyGra – that facilitates scalable, configurable, and high-fidelity generation of synthetic data tailored for these training paradigms. Our approach employs a modular and configuration-based pipeline capable of modeling complex dialogue flows with minimal manual intervention. This framework uses a dual-stage quality tagging mechanism, combining heuristic rules and LLM-based evaluations, to automatically filter and score data extracted from OASST-formatted conversations, ensuring the curation of high-quality dialogue samples. The resulting datasets are structured under a flexible schema, enabling seamless integration into diverse training workflows. Together, these innovations offer a robust solution for generating and managing synthetic conversational data at scale, significantly reducing the overhead of data preparation in LLM training pipelines. Our code and documentation are available at https://github.com/ServiceNow/SyGra
Keywords: Synthetic Data, LLM, DPO, SFT, Graph Pipeline, LangGraph, OASST, Data Generation Framework, Quality Tagging
1 Introduction
The rapid progress of large language models (LLMs) and multimodal AI systems has heightened the demand for large-scale, high-quality training and evaluation datasets [1, 6, 3]. Yet, the cost, bias, and limited availability of annotated real-world data present major barriers [8]. This is especially true in areas like instruction tuning, tool-use supervision, multi-agent interactions, and safety evaluation, where fine-grained control over structure, diversity, and task complexity is essential [4, 5].
Synthetic data, generated via LLMs and automated pipelines, offers greater flexibility and control than traditional datasets. Achieving this at scale, however, poses significant challenges: designing complex, branching workflows that mirror task hierarchies; orchestrating diverse model backends, APIs, and tool calls; enforcing validation and schema compliance across large, heterogeneous outputs; and enabling resumability, sharding, and streaming for scalable, fault-tolerant execution. Reusable, modular flows are also vital for maintainable pipelines.
For teams building domain-specific assistants—such as AI copilots, ticket triaging agents, or safety evaluators—these challenges lead to higher manual effort and slower iteration. A framework is needed that automates high-quality data generation, supports structured outputs and multimodal inputs, and streamlines augmentation—ultimately accelerating the development of custom LLMs for enterprise and research applications.
To address this, we introduce SyGra (Graph-oriented Synthetic-data Pipeline), a general-purpose framework for scalable synthetic data generation. SyGra combines low-code, YAML-based configuration with modular, graph-driven orchestration to support complex workflows with branching, looping, and conditionals. It enables the reuse of graphs as subgraphs, ensures reliable execution through integrated validation and checkpointing, and natively supports multimodal inputs and agent based data generation. Additionally, SyGra offers unified dataset I/O across HuggingFace and local formats, supports quality tagging, and produces outputs compatible with OASST-style formatting for seamless downstream use.
Figure 1: High-level SyGra architecture.
2 Related Work
Recent years have seen rapid progress in the development of synthetic data generation frameworks and instruction-tuning toolkits, with each system making distinct trade-offs across orchestration, extensibility, code abstraction, and multimodal support. Table 1 summarizes core capabilities across representative frameworks like Distilabel [^1] https://distilabel.argilla.io/latest/, SDG [^2] https://github.com/argilla-io/synthetic-data-generator, Curator [^3] https://github.com/bespokelabsai/curator, and Synthetic Data Kit [^4] https://github.com/meta-llama/synthetic-data-kit.
  • Existing data generation frameworks address only subsets of the end-to-end data generation pipeline, leaving gaps in orchestration, extensibility, and multimodal support.
  • Most tools support some combination of asynchronous execution, low-code authoring, configuration-driven flows, and HuggingFace integration, but often lack reusable subgraphs, seamless UI-based workflow design, comprehensive agent/tool support, and integrated quality tagging.
  • UI-based flow configuration is present in some tools (e.g., Curator), but these typically lack robust agent capabilities, multimodal I/O, or subgraphs.
In summary, while existing frameworks each offer valuable features for synthetic data generation, they typically address isolated aspects of the broader workflow. SyGra stands out by providing a unified, extensible approach that brings together the critical capabilities needed for modern, complex, and multimodal data generation pipelines.
3 SyGra Framework
Table 1: Comparison of SyGra with popular frameworks across key capabilities.
Category Feature SyGra Distilabel SDG Curator Synthetic Data Kit
Execution & Authoring Async Execution
Low-Code Authoring
UI-Based Flow Config $\triangle$
Workflow Orchestration Configuration-driven Complex Flow * *
Reusable Subgraphs
Evaluation & Integration Quality Tagging * *
HuggingFace Integration
Agent/Tool Support
Multimodality Multimodal Input * * *
Multimodal Output
  • : Supported    : Not Supported    $\triangle$: Work in Progress    *: Partial Support
SyGra is a modular and extensible system designed for large-scale, programmable data generation. It supports configurable orchestration through a graph abstraction that enables reusable, auditable, and resumable workflows. The framework is designed for both research and production pipelines, with pluggable model backends and modular task authoring support.
3.1 System Architecture
SyGra is guided by three principles—Scalability (streaming data sources, resumable jobs, JSONL/Parquet/HF outputs), Modularity (YAML-defined DAG workflows with conditional logic), and Reusability (versioned, reusable graphs, nodes, and validators). Figure 1 shows its core components:
  1. 1.
    Data I/O: Unified loader/sink for HuggingFace or local CSV, JSON(L), and Parquet in batch or streaming modes.
  2. 2.
    Graph Construction: YAML-defined DAG of nodes (LLM calls, transformations) with conditional edges and pre/post hooks, compiled via LangGraph.
  3. 3.
    Execution Engine: Asynchronous runtime coordinating local Python steps and remote inference (HTTP, OpenAI, Mistral) across VLLM, TGI, OLLAMA and Azure backends, with built-in retries and failure tracing.
  4. 4.
    Structured Output & Resumability: Generates OASST-compatible[2] records and tracks progress metadata for fault-tolerant, restartable runs.
3.2 Pipeline Components
SyGra pipelines are defined declaratively in YAML, promoting low-code, reproducible workflow construction. Each pipeline consists of three configuration blocks:
Data Configuration (data_config).
Specifies input and output sources, format handling (CSV, JSONL, Parquet), streaming options, and inline preprocessing (e.g., renaming, filtering, and combining). Supports both data-backed and data-less generation scenarios.
Graph Configuration (graph_config).
Defines a DAG of computational nodes (Figure 2), each node can be configured to call LLMs, Python functions, agents, or subgraphs. Various node types are supported, such as:
  • llm: When a model needs to be called, we can use a LLM node model properties and role-based prompts, along with pre/post processors in Python code.
  • multi_llm: When we need to generate data at scale, we can use a multi-LLM node which allows configuration of load balanced model inferences between multiple endpoints.
  • lambda: To process the data during execution, we can utilize lambda nodes, which are mapped to Python functions.
  • agent: To perform end-to-end agentic behaviour, we can use agent nodes along with tools, which can be custom or Langchain tools.
  • subgraph: Complex flows can be splitted into smaller graphs i.e. subgraphs which can be reused inside the graph.
Once the nodes are defined, we connect them via egdes (Figure 3). Two types are supported – simple edge and conditional edge. Conditional edges are useful to build if-else flow and loops in the graph based on a condition written as a Python code.
Output Configuration (output_config).
Controls how graph states are serialized into structured output. Users can declaratively map, transform, or customize output using Python hooks to match target schemas like OASST.
Schema Validation.
Ensures output integrity via type and rule-based validation. Schemas can be defined in YAML or Python (e.g., Pydantic), with invalid records automatically skipped and logged.
Finally, the graph is validated and compiled into a LangGraph-compatible representation. Refer to Appendix A for detailed configuration options and schema definitions.
Figure 2: Graph nodes
Figure 3: Graph edges
3.3 Key Features
SyGra brings together robust design abstractions and practical scalability for real-world use cases. Specifically, our contributions include:
  1. 1.
    Low-Code, Modular Graph Configuration: SyGra combines a YAML-based interface with LangGraph-style agents and a custom DAG engine, enabling concise, extensible definitions of complex workflows with branching, looping, and conditionals. B
    data_config:
    source:
    type: "hf"
    repo_id: "google-research-datasets/mbpp"
    config_name: "sanitized"
    split: ["train"]
    graph_config:
    nodes:
    generate_answer:
    node_type: llm
    prompt:
    - system: |
    You are an assistant tasked with solving python coding problems.
    - user: |
    {prompt}
    model:
    name: gpt-4o
    parameters:
    temperature: 0.1
    # more nodes defined here like critique answer
    edges:
    - from: START
    to: generate_answer
    - from: generate_answer
    to: critique_answer
    - from: critique_answer
    to: END
    output_config:
    output_map:
    id:
    from: "id"
    conversation:
    from: "messages"
  2. 2.
    Reusable Recipes (Subgraphs): This feature enables us to use common graph components which can be reused across tasks, promoting modularity. For instance, the Evolve INSTRUCT recipe (Figure  4) encapsulates a modular subgraph that receives seed instructions and applies either depth-based or breadth-based evolution strategies via a routing node (Strategy) [7]. This subgraph can be invoked repeatedly across different flows, enhancing composability and reducing redundancy.
    Figure 4: Instruction evolution subgraph and judgment loop used within SyGra pipelines.
  3. 3.
    Multimodal Support: SyGra extends beyond text-only workflows by natively handling audio and image inputs alongside text. Through unified I/O adapters, it transparently loads local or remote media in various formats, encodes them as base64 data URLs for LLM API compatibility, and supports multiple media fields per record. This enables workflows for tasks such as speech recognition, audio classification, document analysis, and visual QA. Round-tripping ensures outputs can be saved back into HuggingFace datasets in their original formats for reproducibility and downstream use.
    identify_animal:
    output_keys: animal
    node_type: llm
    prompt:
    - user:
    - type: text
    text: |
    Identify the animal in the provided audio.
    - type: audio_url
    audio_url: "{audio}"
[…]
    name: qwen_2_audio_7b
    parameters:
    max_tokens: 1000
    temperature: 0.3
  4. 4.
    Agentic Execution: SyGra enables the creation of autonomous, tool-using agents built on the ReAct reasoning-and-acting paradigm via LangGraph. Agent nodes extend LLM nodes with capabilities for dynamic tool invocation, multi-turn reasoning, and conditional decision-making. Developers can specify a library of callable tools, inject context-specific system messages at arbitrary conversation turns, and configure pre/post-processing hooks for fine-grained control over input and output. This allows pipelines to handle exploratory tasks, iterative search, and interactive decision flows in a modular, low-code manner.
    research_agent:
    node_type: agent
    prompt:
    - system: |
    You are a research assistant that helps users find information.
    Always think step by step and explain your reasoning.
    - user: |
    Please help me research {topic}.
    tools:
    - tasks.sim.tools.search_tool.search
    - tasks.sim.tools.calculator_tool.calculate
    inject_system_messages:
    2: "Remember to cite your sources."
    output_keys:
    - agent_response
    model:
    name: vllm_model
    parameters:
    temperature: 0.2
    max_tokens: 1024
  5. 5.
    Structured Output Generation: SyGra provides a flexible framework for generating and validating structured outputs from LLMs, reducing post-processing effort and ensuring reliable formats. It supports both class-based schemas (via Pydantic) and YAML-defined schemas, with automatic type handling and optional custom validation rules. Structured output generation works natively with OpenAI and vLLM models, and falls back to JSON schema validation for other backends. This allows developers to define precise field types, attach descriptions, and enforce constraints directly at generation time.
    nodes:
    answer_node:
    node_type: llm
    model:
    name: gpt-4o
    parameters:
    temperature: 0.1
    structured_output:
    enabled: true
    schema:
    fields:
    answer:
    type: str
    description: "Main answer text"
    confidence:
    type: float
    description: "Confidence score between 0 and 1"
  6. 6.
    Resumability: SyGra supports fault-tolerant, restartable execution of long-running jobs. In the event of a failure, execution can gracefully shut down and later resume from the last recorded checkpoint without reprocessing completed steps. This is particularly valuable for large-scale or streaming workloads where partial progress should be preserved. Checkpoints store both intermediate outputs and node-level metadata, enabling accurate restoration of execution state.
    python main.py task <your_task> resume True
  7. 7.
    Filterable OASST-Compatible Formatting: Outputs can be structured in an OASST-compatible [2] format for easy post-hoc filtering, inspection, and training integration.
    Figure 5: An example Conversation Tree of depth 4 containing 12 messages [2]
4 Results and Impact
4.1 Experimental Setup
The evaluation was run on an 8-core CPU machine with 16 GB RAM using the SyGra framework. Model endpoints were deployed separately on vLLM with the Qwen 3 32B Instruct model, configured to use two GPUs with tensor parallelism, moderate CPU and memory resources, and optimizations such as chunked prefill with capped GPU memory usage.
The workload consisted of 10,000 input records, repeated across three trials at each concurrency level. To keep the focus on system behavior, the workflow was deliberately simple: a weighted sampler node selected tone and persona values, and an LLM node rephrased the input text using fixed generation parameters (deterministic outputs, max length 500 tokens). This design provided a controlled inference workload for measuring concurrency scaling without pipeline complexity.
4.2 Results
Performance was measured in terms of total wall-clock time required to process the workload under varying concurrency levels (denoted here as batch size).
Figure 6: Total completion time (seconds) for varying concurrency levels with one endpoint.
Three key patterns emerge:
  • Reduced completion time with higher concurrency. As concurrency increases from 10 to $\sim$500, total completion time decreases sharply, illustrating the efficiency of SyGra’s asynchronous execution.
  • Sustained efficiency at scale. Beyond $\sim$500 concurrent requests, completion time stabilizes around 900–1000 seconds, showing resilience under heavy load.
  • Server saturation at very high concurrency. At batch size 5000, total time increases slightly. This reflects a limitation of the underlying model server rather than SyGra itself. Users can mitigate this by adding multiple endpoints or increasing the computational resources allocated to the hosted model.
Note on Task Complexity. The experiment used a deliberately simple two-node workflow (sampling + rephrasing) to isolate concurrency effects. In this setting, a single vLLM endpoint was sufficient to handle high concurrency. However, in more complex tasks with multiple nodes and richer pipelines, SyGra exhibits stronger scaling when additional endpoints or resources are available, as state-of-the-art inference servers such as vLLM are capable of sustaining concurrent requests efficiently across distributed deployments.
4.3 Impact
These results underscore SyGra’s ability to deliver stable, predictable performance under extreme concurrency. Whereas traditional multi-threaded or multi-process inference servers often degrade sharply once concurrency exceeds a few hundred requests, SyGra sustains bounded completion times even at >>5000 concurrent requests.
The implications are significant:
  • Scalable deployments. SyGra enables operators to support thousands of simultaneous queries while maintaining consistent workload completion times.
  • Configurable performance. If saturation is observed at very high concurrency, performance can be extended by scaling to multiple endpoints or allocating additional resources to the hosted model.
  • Future readiness. As LLM-based systems (e.g., conversational agents, retrieval-augmented generation) face growing concurrency demands, SyGra provides a reliable inference layer that avoids bottlenecks at production scale.
In summary, SyGra demonstrates robust concurrency scaling, endpoint-agnostic performance, and stable execution times, making it a strong candidate for next-generation large-scale data syntheses.
5 Availability
SyGra is released as an open-source Python package and framework on PyPI555pip package: https://pypi.org/project/sygra/ (pip install sygra) and GitHub [^6] https://github.com/ServiceNow/SyGra.
6 Conclusion
We presented SyGra, a modular framework for synthetic data generation using graph-based, prompt-centric workflows. SyGra offers scalable, reproducible pipelines for language model training, featuring a low-code YAML interface, reusable subgraphs, agent nodes, and HuggingFace-native I/O. Its design supports diverse workflows, uniquely enabling multimodal inputs, subgraph reuse, conditional routing, and schema validation.
Current limitations include text-only outputs (with multimodal output support planned), independent node operation without cross-sample reasoning, and basic agent support. Future work will address richer agent coordination, stateful nodes, real-time graph editing, and uncertainty modeling.
SyGra accelerates dataset creation and promotes transparency and reuse in LLM development. Ongoing efforts must address risks like “model collapse” through mixed datasets and continuous quality control, ensuring SyGra’s utility across generative AI applications.
Appendix A Pipeline Components: Features, Definitions and Examples
A.1 Data Configuration
A.1.1 Input Sources
Sample YAML Configuration:
Figure 7: An example configuration using a HuggingFace dataset as source and applying field renaming transformation is shown below.
This configuration illustrated in Figure 7 represents:
  • Input from HuggingFace and local disk (alternative)
  • Use of RenameFieldsTransform for renaming schema fields
  • Optional sink setup with HuggingFace or local file export
A.1.2 Transformations
RenameFieldsTransform.
The RenameFieldsTransform is a lightweight transformation utility used in the SyGra pipeline to rename one or more fields in each record of the dataset. This is particularly useful for ensuring consistency in variable naming, aligning raw data to prompt-ready formats, or preparing input fields for downstream processing.
The YAML configuration for this transformation accepts a mapping parameter, which specifies how input field names should be renamed. An optional overwrite flag determines whether to overwrite any existing field in case of name collision.
Figure 8(a) shows a sample usage where the fields page, llm_extract, and type are renamed to id, text, and text_format, respectively.
(a) Example usage of RenameFieldsTransform in YAML configuration. This renames selected fields to align with graph input expectations.
(b) CombineRecords configuration: aggregates two records with shift and join logic.
(c) SkipRecords configuration: skips first and last 10 records using count-based range.
(d) Minimal YAML configuration demonstrating data-less mode with only a sink field.
Figure 8: Examples of YAML configurations used for field renaming, record combination/skipping, and minimal data sink setup.
CombineRecords.
This transformation combines multiple records to form richer contextual input. It can skip from the beginning or end of the dataset, define how many records to combine, and how to shift the combination window. As shown in Figure 8(b), the configuration merges two records, joining multiple fields with newline delimiters or preserving the first record’s values.
SkipRecords.
Figure 8(c) presents a simpler configuration to exclude records from the dataset, either from the start or end. This is especially useful for filtering noisy, incomplete, or structurally incompatible entries prior to processing.
A.1.3 Data Less Mode
In data-less mode, SyGra operates without any input source. Instead, it directly executes the graph and writes outputs based solely on intermediate or generated values. This is especially useful for bootstrapping datasets, performing zero-shot synthesis, or generating instructional data.
Figure 8(d) shows a minimal YAML configuration that defines only an output sink.
A.2 graph_config: Nodes and Execution Flow
Graph-Level Properties:
  • chat_conversation: singleturn or multiturn
  • chat_history_window_size: integer
Node Types:
  • llm — standard prompt inference
  • multi_llm — ensemble-style multi-model generation
  • weighted_sampler — controlled randomness
  • lambda — run Python logic
  • agent — multi-turn agent execution with memory and tools
  • subgraph — reusable logical block
Each node can define:
  • Prompt templates with variable substitution
  • Model name and parameters
  • Input/output keys, chat history, role labeling
  • Pre-process and post-process functions
Edge Types:
  • Simple Edges: Direct transitions between nodes.
  • Conditional Edges: Conditional routing via Python classes and path_map.
Special nodes: START and END are implicit entry and exit points.
A.3 output_config: Record Generation
Declarative Output Mapping: Each field in output_map can use:
  • from: Reference a graph state variable
  • value: Assign a static constant
  • transform: Apply method in generator class
Supports context-aware templating with \$ paths to inject YAML metadata (e.g., \$data_config.source.repo_id).
Custom Output Generators: Advanced logic can override the generate() method to control formatting or field post-processing.
A.4 schema_config: Output Validation
SyGra supports both declarative and programmatic schema enforcement:
Option 1: YAML-based Schema
  • Define fields with name, type, and optional rules (e.g., is_greater_than, regex).
Option 2: Python Schema Class
  • Define a class extending BaseModel, use Pydantic @validator or @root_validator.
Validation is applied post-execution; failing records are logged and skipped.
# Example A: use a custom Pydantic schema class
schema_config:
schema: validators.custom_schemas.CustomUserSchema
# Example B: inline field schema with rules
schema_config:
fields:
- name: id
type: int
is_greater_than: 99999 # ensure >= 6 digits
- name: conversation
type: list[dict[str, any]]
- name: taxonomy
type: list[dict[str, any]]
- name: annotation_type
type: list[str]
- name: language
type: list[str]
- name: tags
type: list[str]
A.5 Post-Generation Extensions
OASST Mapper: Enables conversion of records into SFT/DPO format based on the OpenAssistant schema. Activate with: --oasst True
Quality Tagging: Automatically tags records using LLMs or heuristics. Enable with: --quality True
Appendix B Example SyGra YAML Configurations
This appendix provides example YAML configurations illustrating how SyGra pipelines are defined and composed using the data_config, graph_config, output_config, and schema_config sections. These examples demonstrate SyGra’s flexibility for data-driven and zero-shot pipelines, LLM orchestration, and safe output generation.
B.1 Minimal Data-Less Generation Configuration
data_config:
sink:
type: "json"
file_path: "output/synthetic_data.jsonl"
[…]
nodes:
generate:
node_type: llm
output_keys: response
prompt:
- system: "You are a helpful assistant."
- user: "Write a fun fact about space."
model:
name: gpt-3.5-turbo
parameters:
temperature: 0.8
edges:
- from: START
to: generate
- from: generate
to: END
[…]
output_map:
fact:
from: response
B.2 Full Pipeline with Conditional Edge and Schema Validation
data_config:
source:
type: "disk"
file_path: "data/code_tasks.jsonl"
file_format: "jsonl"
sink:
type: "jsonl"
file_path: "output/validated_output.jsonl"
[…]
nodes:
generate:
node_type: llm
output_keys: solution
prompt:
- system: "You are an AI that solves code problems."
- user: "{task}"
model:
name: mistral
parameters:
temperature: 0.5
validate:
node_type: lambda
lambda: validators.code.check_validity
output_keys:
- is_valid
edges:
- from: START
to: generate
- from: generate
to: validate
- from: validate
condition: validators.code.RouteBasedOnValidity
path_map:
END: END
generate: generate
[…]
output_map:
id:
from: task_id
solution:
from: solution
validity:
from: is_valid
[…]
fields:
- name: id
type: int
- name: solution
type: str
- name: validity
type: bool
B.3 Pipeline to process images as input
data_config:
source:
type: "hf"
repo_id: "datasets-examples/doc-image-1"
split: "train"
streaming: true
[…]
type: "hf"
repo_id: <repo_name>
config_name: $M$-doc-image-1
split: train
push_to_hub: true
private: true
token: <hf_token>
[…]
nodes:
judge_pokemon:
output_keys: pokemon
node_type: llm
prompt:
- user:
- type: text
text: |
Identify the pokemon in the provided image.
- type: image_url
image_url: "{image}"
[…]
name: gpt-4o
parameters:
max_tokens: 1000
temperature: 0.3
edges:
- from: START
to: judge_pokemon
- from: judge_pokemon
to: END
[…]
output_map:
id:
from: "id"
image:
from: "image"
pokemon:
from: "pokemon"
[…]
B.4 Pipeline to process audio inputs
data_config:
source:
type: "hf"
repo_id: "datasets-examples/doc-audio-1"
split: "train"
streaming: true
[…]
nodes:
identify_animal:
output_keys: animal
node_type: llm
prompt:
- user:
- type: text
text: |
Identify the animal in the provided audio.
- type: audio_url
audio_url: "{audio}"
[…]
name: qwen_2_audio_7b
parameters:
max_tokens: 1000
temperature: 0.3
edges:
- from: START
to: identify_animal
- from: identify_animal
to: END
[…]
output_map:
id:
from: "id"
audio:
from: "audio"
animal:
from: "animal"
[…]